OpenStudy (anonymous):

solve the diff eq dy\dx+3x^2*y=x^5*e^x^3

OpenStudy (anonymous):

this follows the same method

OpenStudy (anonymous):

as above

OpenStudy (anonymous):

Yes, Here u Chose U as U = e^int( 3x^2)

OpenStudy (anonymous):

\[U \times y \int\limits_{}^{} U \times x ^{5\times}e ^{x ^{3}}\]

OpenStudy (anonymous):

no...not required, u can take u=x^3

OpenStudy (anonymous):

aha, count it as a seperabel?

OpenStudy (anonymous):

after simplification, the eq. becomes \[y dy = (1/3) (x^3 . e^ (x^3))

OpenStudy (anonymous):

ignore \[ in the beginning

OpenStudy (anonymous):

so in RHS, u can take u=x^3 and solve....okay?

OpenStudy (anonymous):

Are u sure u have done right ..

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

aren't you getting the same?

OpenStudy (anonymous):

noo?

OpenStudy (anonymous):

\[Dy \div Dx + 3x ^{2} y=x ^{5} e ^{x ^{3}}\]

OpenStudy (anonymous):

he means i hope..

OpenStudy (anonymous):

yeah its the same eq the ans is \[12y=e ^{x ^{3}}(2x^3-1)+ce ^{-x ^{3}}\] how to get that?